OpenStudy (kingj):

The slope of a line is 2/3 What is the slope of a line that is perpendicular to it? -3/2 -2/3 3/2 3/2

OpenStudy (campbell_st):

if the slope of the line is 2/3, then the perpendicular is the negative reciprocal \[m _{1} = \frac{2}{3} .... m _{2} = -\frac{3}{2}\] the slopes must follow the rule \[m _{1} \times m _{2} = -1\]